### v1

# ESIQA: Perceptual Quality Assessment of Vision-Pro-based Egocentric Spatial Images

Abstract
With the development of eXtended Reality (XR), head-mounted shooting and display technology have experienced significant advancement and gained considerable attention. Egocentric spatial images and videos are emerging as a compelling form of stereoscopic XR content. Different from traditional 2D images, egocentric spatial images present challenges for perceptual quality assessment due to their special shooting, processing methods, and stereoscopic characteristics. However, the corresponding image quality assessment (IQA) research for egocentric spatial images is still lacking. In this paper, we establish the Egocentric Spatial Images Quality Assessment Database (ESIQAD), the first IQA database dedicated for egocentric spatial images as far as we know. Our ESIQAD includes 500 egocentric spatial images, containing 400 images captured with the Apple Vision Pro and 100 images generated via an iPhone’s “Spatial Camera” app. The corresponding mean opinion scores (MOSs) are collected under three viewing modes, including 2D display, 3D-window display, and 3D-immersive display. Furthermore, based on our database, we conduct a benchmark experiment and evaluate the performance of 22 state-of-the-art IQA models under three different viewing modes. We hope this research can facilitate future IQA research on egocentric spatial images. The database is available at https://github.com/IntMeGroup/ESIQA.
1 Introduction
Egocentric shooting and displaying technology has gained considerable attention recently, which can provide convenient capturing methods and immersive experiences through head-mounted displays (HMDs) [7, 4, 5], such as Apple Vision Pro, Meta Quest 3, and Google Glass, etc. Egocentric spatial images, captured from the first-person perspective, are emerging as an important form of stereoscopic content. As shown in Figure 1, egocentric spatial images are captured with HMDs in real-world scenes and can be transmitted and viewed on various terminal devices. For binocular displays, since a spatial image is composed of a pair of disparity views to simulate human binocular vision, it is possible to directly create a three-dimensional (3D) effect and provide an immersive experience. For conventional flat displays, generally only the left view, i.e., a standard two-dimensional (2D) image, is presented.
Figure 1: Illustration of the capturing and displaying methods for egocentric spatial images.
Egocentric spatial images, taken by a wide variety of individuals from professional photographers to amateurs, exhibit significant differences in visual quality [7, 19]. This quality variation is further compounded by the dynamic and uncontrolled environments when capturing egocentric spatial images, leading to common distortions such as under/over exposure, low visibility, noise, color shift, etc [4]. Moreover, since the captured spatial images can be viewed on any device, the corresponding quality of experience (QoE) may also be different. Therefore, it is important to develop image quality assessment (IQA) metrics of egocentric spatial images to better maintain and optimize QoE on various devices.
Figure 2: Sample egocentric spatial images from our ESIQAD.
Over the past few decades, many studies have focused on the traditional 2D image quality assessment problem, with the establishment of several IQA databases, e.g., TID2013 [23], KonIQ-10K [15], and KADID-10k [16], and the development of many advanced IQA models [21, 18, 35, 31]. However, traditional IQA models are inadequate for assessing the quality of egocentric spatial images due to their unique immersive experiences, which have a distinct impact on QoE. Moreover, spatial images consist of two distinct views (left and right), which precludes the direct application of traditional IQA models. With the diversification of image forms, some studies have focused on 3D IQA (also known as stereoscopic IQA) [22, 10, 8, 36] and VR IQA [28, 3, 37, 6]. However, existing 3D IQA metrics fail to consider the immersive experience of egocentric spatial images, and VR IQA metrics cannot effectively evaluate their stereoscopic effects. These metrics are unsuitable for assessing the quality of egocentric spatial images. As far as we know, there is a lack of IQA studies for egocentric spatial images, which have significant differences compared with traditional 2D images, 3D images, and VR images in terms of formats, characteristics, and applications. Therefore, there is an urgent need for dedicated research on egocentric spatial image quality assessment.
To better understand human visual preferences for egocentric spatial images and facilitate the development of IQA models tailored to these images, we construct the first Egocentric Spatial Images Quality Assessment Database, termed ESIQAD. The database contains 500 egocentric spatial images with diverse scenes and the corresponding collected human perceptual quality ratings on three viewing modes, including 2D, 3D-window, and 3D-immersive. Specifically, 400 spatial images are captured using the Apple Vision Pro, and 100 spatial images are generated by the images taken by the iPhone’s “Spatial Camera” application. Based on the collected egocentric spatial images, we conducted subjective experiments to evaluate the QoE of these images, in which 22 participants were included to provide their perceptual quality ratings under three viewing modes. Based on our ESIQAD, we conduct a benchmark experiment to evaluate the performance of 22 state-of-the-art IQA models under three viewing modes. Our contributions are summarized as follows:
  • We establish a large-scale quality assessment database for egocentric images, named ESIQAD, which is the first IQA database for egocentric spatial images to the best of our knowledge.
  • We analyze the human preference characteristics for egocentric spatial images under three viewing modes based on our ESIQAD.
  • We conduct a benchmark experiment by evaluating the performance of numerous state-of-the-art IQA models based on our ESIQAD.
2 Database Construction
To facilitate the egocentric spatial IQA research, we construct a large-scale database, termed ESIQAD, and conduct the corresponding subjective experiment to derive the subjective quality scores in terms of three viewing modes.
Figure 3: Three viewing modes in our subjective experiments.
2.1 Content Collection
Firstly, we utilized the Apple Vision Pro and the iPhone 12 as the shooting devices to collect spatial images from a variety of scenes, including both indoor and outdoor environments, different weather conditions, and various times of day. These images feature various characteristics including brightness, contrast, colorfulness, spatial details, etc., and have different degradations including noise, blur, etc. The raw images were taken by ten photographers, representing a wide range of user perspectives and preferences. Specifically, we collected 400 spatial images captured with the Apple Vision Pro, and 100 spatial images generated from the images taken by the iPhone, referred to as captured images and generated images, respectively. For the 100 generated images, we used the “Spatial Camera” app to automatically take a 2D image as the left view and then generate the right view to form a spatial image. It should be noted that the 100 generated images have their corresponding images with the same scenes in the 400 images captured by Apple Vision Pro to facilitate the study of human visual preferences. The captured images have a fixed resolution of 2560 × 2560 per view, while the generated images have a resolution of 4032 × 3024 per view. Figure 2 demonstrates sample images from our ESIQAD, where all samples are illustrated in the main view (left view).
2.2 Subjective Experiment
2.2.1 Experiment setup
Based on the collected images, we then conducted a subjective experiment to gather quality scores of human perception on egocentric spatial images across different viewing modes, including 2D, 3D-window, and 3D-immersive, as shown in Figure 3. For the 2D viewing mode, egocentric spatial images were displayed on a 23.8-inch Redmi T238NF monitor with the resolution of 1920 × 1080, driven by a Mac Mini M2 PC. We developed a quality scoring interface using Python Tkinter, which was used to present egocentric spatial images (only the left view) and collect subjective quality ratings. The interface allows users to freely browse the previous and next images, and record the quality scores through a scroll bar. For the 3D-window and 3D-immersive viewing modes, we utilized the Apple Vision Pro device as the head-mounted display, which features the resolution of 3660 × 3200 per eye, the 120°$\degree$ horizontal field of view (FoV), and the 100Hz refresh rate. For the 3D-window mode, egocentric spatial images appeared as small windows within the HMD, allowing users to maintain awareness of the surroundings. For the 3D-immersive mode, egocentric spatial images were shown in a wider immersive FoV, which provides an immersive egocentric experience. In both 3D viewing modes, all images were displayed through the built-in application “Photos” and displayed at their native resolutions to prevent the scaling distortion. Across all three viewing modes, egocentric spatial images were presented in a random order.
2.2.2 Experiment methodology
We adopted the single stimulus absolute category rating (SSACR) method to collect subjective quality ratings for egocentric spatial images under three viewing modes, with the rating scale ranging from 1 to 10. Under the 2D viewing mode, the viewers were seated at a distance of about 2 feet from the monitor in a laboratory environment with normal indoor illumination, and utilized the scoring interface to assign quality scores. Under the two 3D viewing modes, participants were instructed to wear the Apple Vision Pro and select either the 3D-window or 3D-immersive viewing mode via the “Photos” app. A keyboard was placed in front of the participants to enter the scores. Participants manually entered their scores after viewing each image and proceeded to the next image using gestures.
A total of 22 graduate students (11 males and 11 females) were invited to participate in the subjective experiment. Before the experiment, all participants completed a standard procedure in visual testing, i.e., the Snellen visual acuity test [1]. All participants were confirmed normal or corrected-to-normal vision. For all three viewing modes, the participants were instructed to assess the egocentric spatial images based on both distortion and aesthetic quality, providing ratings with an overall quality score. The whole experiment is split into 2 sessions, each of which includes a subjective quality evaluation process for 750 samples (250 spatial images $\times$ 3 modes). All participants completed both sessions, producing a total of 33,000 quality ratings (22 participants $\times$ 1,500 ratings). After the scoring sessions, each participant was asked to answer a questionnaire regarding their experience. Details of the questionnaire are given in Section 2.3.2.
Figure 4: Distribution of MOSs differences of spatial images across three modes.
2.3 Subjective Data Processing and Analysis
2.3.1 MOS Calculation and Analysis
We perform the same processing methods for the collected scores under three viewing modes respectively. Firstly, we follow the subjective data processing guidelines recommended by ITU [24] for outlier detection and subject rejection. For 2D and 3D-immersive viewing modes, 1 of the 22 participants was deemed as an outlier and excluded, respectively. For the 3D-window viewing mode, none of the 22 participants were excluded. Then, we convert the raw scores from the subjects into normalized Z-scores, which range between 0 and 100. Subsequently, we calculate the average of these Z-scores to derive the mean opinion scores (MOSs), which are formulated as follows:
$z_{ij}=\frac{r_{ij}-\mu_{i}}{\sigma_{i}},\quad z_{ij}^{\prime}=\frac{100\left(z_{ij}+3\right)}{6},$ (1)
$\text{MOS}_{j}=\frac{1}{N}\sum_{i=1}^{N}z_{ij}^{\prime},\vspace{-4pt}$ (2)
where $r_{ij}$ is the original score of the $i$-th subject on the $j$-th image, $\mu_{i}$ and $\sigma_{i}$ represent the mean rating and the standard deviation given by subject $i$, respectively, and $N$ is the total number of subjects.
Figure 5: Distribution of MOSs differences of spatial images across three modes. (a) $\text{MOS}_{\text{3D-window}}-\text{MOS}_{\text{3D-immersive}}$. (b) $\text{MOS}_{\text{3D-immersive}}-\text{MOS}_{\text{2D}}$. (c) $\text{MOS}_{\text{3D-window}}-\text{MOS}_{\text{2D}}$.
In the following analysis, we utilize original MOSs, ranging from 1 to 10, instead of normalized scores to facilitate clearer comparisons across different viewing modes. Figure 4 (a) presents the histograms of the MOS distribution over the three viewing modes, indicating that the perceptual quality scores are widely distributed in the $\left[1,10\right]$ interval in each mode. In addition, Figure 4 (b) and (c) demonstrate the MOS distributions of 400 captured images and 100 generated images, respectively. It can be observed that the MOS distribution for captured images centers around 5 in all three modes, while it centers above 5 for generated images. This is mainly due to the fact that the color of the images captured by the iPhone is generally enhanced to be more comfortable for human vision perception, while the color of the images captured by Apple Vision Pro is generally kept real. Moreover, for captured images, the perceptual quality is higher in 3D-immersive and 3D-window viewing modes compared to 2D viewing mode. However, for generated images, the perceptual quality is better in 2D viewing mode than in 3D viewing modes.
The results shown in Figure 4 (a) indicate that the MOS fitting curves for the two 3D viewing modes are located at the right of the 2D viewing mode, manifesting higher perceived image quality in the 3D environments. To facilitate a more direct comparison of the perceptual quality of the same image across different viewing modes, we calculate the MOS difference for each image between the three modes. The differences are illustrated in the frequency distribution histograms as shown in Figure 5. The results indicate that the perceptual quality of most images in 3D-window and 3D-immersive scenes is comparable but generally superior to that observed on 2D displays.
Figure 6: Distributions of the MOS difference ($\text{MOS}_{\text{generated}}-\text{MOS}_{\text{captured}}$) between 100 matched pairs of captured and generated images under three viewing modes.
To compare the perceptual quality of captured and generated images with identical scenes, we compute the MOS difference between 100 matched pairs of captured and generated images across three viewing modes by subtracting the MOS of generated images and their corresponding captured images. The distribution histogram of the differences is depicted in Figure 6. The results indicate that the perceptual quality of most generated images exceeds that of their corresponding captured images. This improvement mainly results from image enhancement methods applied during the image generation.
Figure 7: A list of the questions included in the post-test questionnaire.
2.3.2 Questionnaire Analysis
Post-test questionnaires can help better understand the QoE of egocentric spatial images and design future studies. All questions are shown in Figure 7. For Q1 and Q2, we derive ranking scores for the three viewing modes based on the sorting results, which reflect the overall ranking of the options. The ranking score for a viewing mode is calculated by
$S=\frac{\sum(f\times w)}{n},$ (3)
where $S$ indicates the ranking score for this viewing mode, $f$ represents the frequency of each rank, $w$ denotes the weight assigned to each rank (e.g., the top rank is weighted 3, and the lowest rank is weighted 1), and $n$ is the total number of participants. For Q3 and Q4, we compute the average scores from questionnaire responses to evaluate user dizziness. The statistical results of the replies and some general observations are as follows.
  • Q1: 3D-immersive mode, 3D-window mode, and 2D mode obtain ranking scores of 2.27, 2.18, and 1.55, respectively. The results suggest that most subjects prefer the experience of 3D-immersive and 3D-window modes viewing in Apple Vision Pro compared to the experience of the 2D mode viewing on the screen display, which is consistent with our previous analysis in Section 2.3.1.
  • Q2: We derive the ranking scores for each viewing mode: 3.00, 1.73, and 1.14 for 3D-immersive, 3D-window, and 2D modes, respectively. The results indicate participants generally consider that the 3D-immersive scene offers the most immersive experience.
  • Q3 & Q4: The scores for dizziness in the 3D-immersive and 3D-window viewing modes are 4.27 and 2.27, respectively. This result suggests that while the 3D-immersive mode offers a superior immersive experience, it also induces more significant dizziness compared to the 3D-window mode.
2.3.3 Database Attributes Analysis
Following the methodologies outlined in [29] and [12], we conduct statistical analysis for our ESIQAD in terms of five low-level vision feature dimensions, including sharpness, brightness, colorfulness, contrast, and cumulative probability of blur detection metric (CPBD). Considering that the ranges of these features are different, we normalized their values to $\left[0,1\right]$ for better illustration. Figure 8 shows the kernel distribution of each feature for our ESIQAD. It can be observed that the spatial images in our database have a wide distribution in brightness, colorfulness, contrast, and CPBD features, showing their extensive diversity. However, our database shows a narrow distribution range for the sharpness feature, notably around 0.3.
Figure 8: Distributions of five low-level vision features in our ESIQAD.
3 Experiment
Table 1: Performance comparison of the state-of-the-art NR IQA models on our ESIQAD under three viewing modes. The best performances are marked in RED, and the second-best performances are marked in BLUE.
Dimension 2D 3D-window 3D-immersive
Model SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC
BMPRI [18] 0.2803 0.1798 0.5525 0.1255 0.0825 0.2841 0.0850 0.0556 0.2780
BPRI [17] 0.1780 0.1127 0.2422 0.0925 0.0637 0.1380 0.0779 0.0535 0.1400
BPRI-LSSn [17] 0.5800 0.4180 0.6255 0.4287 0.3028 0.5768 0.3550 0.2472 0.4703
BPRI-LSSs [17] 0.1400 0.0746 0.3675 0.1036 0.0741 0.1499 0.1189 0.0823 0.1675
BPRI-PSS [17] 0.2115 0.1446 0.4519 0.2348 0.1624 0.4153 0.1923 0.1320 0.2749
BRISQUE [20] 0.5594 0.3972 0.6642 0.3156 0.2186 0.4960 0.2858 0.1943 0.4623
CORNIA [34] 0.4073 0.2837 0.5134 0.1768 0.1188 0.3515 0.1572 0.1043 0.3555
FISBLIM [11] 0.4543 0.3191 0.5767 0.2800 0.1943 0.4888 0.2238 0.1521 0.4743
HOSA [31] 0.5076 0.3599 0.6211 0.2713 0.1849 0.4218 0.2749 0.1864 0.4298
NIQE [21] 0.5665 0.4088 0.6881 0.3081 0.2106 0.4957 0.3133 0.2137 0.4977
ILNIQE [35] 0.3348 0.2317 0.4187 0.3118 0.2144 0.4113 0.2670 0.1827 0.3702
LPSI [30] 0.5710 0.4071 0.6959 0.3986 0.2786 0.5384 0.3396 0.2349 0.4920
QAC [32] 0.4232 0.2867 0.5638 0.2796 0.1884 0.3882 0.2005 0.1346 0.3881
CNNIQA [14] 0.5824 0.4238 0.5404 0.5517 0.4001 0.4738 0.4579 0.3224 0.4050
Resnet18 [13] 0.7574 0.5628 0.7006 0.6120 0.4554 0.6678 0.5127 0.3754 0.5974
Resnet34 [13] 0.6882 0.5107 0.6632 0.6374 0.4839 0.6492 0.6213 0.4582 0.6952
VGG16 [26] 0.6641 0.4747 0.7031 0.5958 0.4436 0.6718 0.5041 0.3666 0.6125
VGG19 [26] 0.6883 0.5022 0.6916 0.6015 0.4470 0.6747 0.5106 0.3731 0.6007
HyperIQA [27] 0.6925 0.5188 0.6870 0.4901 0.3491 0.4675 0.4013 0.2808 0.4369
MANIQA [33] 0.8532 0.6727 0.8934 0.7894 0.6101 0.8565 0.7539 0.5646 0.7924
TReS [9] 0.7400 0.5620 0.8223 0.6405 0.4696 0.6974 0.6311 0.4525 0.6030
WaDIQaM [2] 0.6598 0.4814 0.6963 0.6493 0.4688 0.6794 0.5010 0.3539 0.5702
3.1 Benchmark Models
To evaluate the performance of various existing IQA models on predicting human visual perception for egocentric spatial images, we employ 22 state-of-the-art no-reference (NR) IQA models for comparison, which are designed for traditional 2D images. The selected models can be categorized into two groups:
  • $\bullet$
    Hand-crafted models, which include BRISQUE [20], CORNIA [34], QAC [32], BMPRI [18], NIQE [21], ILNIQE [35], HOSA [31], FISBLIM [11], LPSI [30], BPRI-PSS [17], BPRI-LSSs [17], BPRI-LSSn [17], and BPRI [17].
  • $\bullet$
    Deep learning-based models, including CNNIQA [14], Resnet18 [13], Resnet34 [13], VGG16 [26], VGG19 [26], HyperIQA [27], WaDIQaM [2], MANIQA [33], and TReS [9].
For traditional hand-crafted IQA models, we apply these models to directly predict the quality scores of the left and right views of an egocentric spatial image, respectively. The predicted score from the left view, which is displayed in the 2D viewing mode, directly serves as the 2D quality score for the image. For both 3D-window and 3D-immersive modes, the quality score for the egocentric spatial image is calculated by averaging the scores from the left and right views.
For deep learning-based IQA models, we train and test these models on our ESIQAD. We partition the database into training and testing sets with the ratio of 4:1. The training parameters are set the same as those in the officially released version. Considering that these deep learning-based IQA models are designed for traditional 2D images, we feed the left view of spatial images into the models to predict quality scores for the 2D viewing mode. Since spatial images cannot be directly input into these models, we fine-tune the architectures of these deep learning-based models to accommodate a dual-image (6 channels) input, comprising both left and right views of the spatial image. These fine-tuned models are then utilized to predict quality scores under 3D-window and 3D-immersive viewing modes.
3.2 Performance Analysis
We evaluate the performance of these benchmark models using three evaluation criteria, i.e., Spearman rank correlation coefficient (SRCC), Pearson linear correlation coefficient (PLCC), and Kendall’s rank correlation coefficient (KRCC). These three metrics are utilized to measure the prediction monotonicity. Before computing these criteria, predicted scores are normalized using a five-parameter logistic function [25]:
$\hat{y}=\beta_{1}\left(0.5-\frac{1}{1+e^{\beta_{2}\left(y-\beta_{3}\right)}}\right)+\beta_{4}y+\beta_{5},$ (4)
where $\left\{\beta_{i}\mid i=1,2,\ldots,5\right\}$ are parameters to be fitted, $y$ and $\hat{y}$ indicate the predicted scores and the corresponding mapped scores, respectively.
Table 1 shows the performance of the aforementioned 22 state-of-the-art models. It can be observed that most hand-crafted models show poor performance in evaluating the perceptual quality of egocentric spatial images in 2D viewing mode, and even perform worse for the evaluation in 3D-window and 3D-immersive modes. Compared to the hand-crafted models, deep learning-based models generally exhibit better performance across all three modes. Among all the IQA models, MAINIQA [33] achieves the best performance across three modes, demonstrating excellent prediction ability and strong generalization capabilities in different modes. Across all models, the performance for 2D-viewing-mode quality assessment is generally better than that for the 3D-window and 3D-immersive settings. The result stems from the fact that these metrics are initially designed for 2D IQA and fail to consider the stereoscopic and immersive effects inherent in 3D scenes, which significantly influence the perceptual quality. Additionally, most models exhibit slightly improved performance in predicting quality in 3D-window modes compared to 3D-immersive modes, which manifests that 3D-immersive mode presents greater challenges in predicting its QoE. In future work, we will utilize the left and right views and their disparities in spatial images to build a quality assessment model with better prediction performance and suitable for different viewing modes.
4 Conclusion
In this work, we establish the first egocentric spatial images database, termed ESIQAD, comprising 400 captured images and 100 generated images along with their MOSs under three viewing modes, i.e., 2D, 3D-window, and 3D-immersive modes. Based on our ESIQAD, we analyze the human visual perception of these images across various viewing modes and conduct a benchmark study by evaluating the performance of 22 state-of-the-art NR IQA models. The results from our benchmark study indicate that most current advanced IQA models fail to accurately assess the perceptual quality of spatial images, particularly under two 3D settings. It is worthwhile to explore the stereoscopic features of spatial images and effectively utilize disparity information to further improve the performance of egocentric spatial image quality assessment metrics in the future.